An emf is induced in a conducting loop of wire 1.38 m long as its shape is changed from square to circular.

Find the average magnitude of the induced emf if the change in shape occurs in 0.165 s and the local 0.478-T magnetic field is perpendicular to the plane of the loop.
Answer #1

emf induced in a conducting loop of wire length I = 1.38m. magnetic field (B) = 0.478T. Shape of cloop changes from square toe =B (T1² - 0²) 0.165 e = 0.478 ( 3.14x (0.878} - (1.38² ) 0.165 e = 2.89 (21420 - 1.904) e = 2.89 (0.516) Je = 11491 emf ind